# Global thermodynamic manifold for conservative control of stochastic systems

### v2

@@ V.1 The Drazin inverse
Intuitively, the need for an appropriately defined inverse of the generator arises due to the time integral in the definition (44) of the full-control friction tensor. In the same way as the generator $\mathcal{G}_{t}$ plays the role of a time derivative on the space of probability distributions, its Drazin inverse $\mathcal{G}^{\mathcal{D}}_{t}$ plays the role of a time integral (over $[0,\infty)$) on the same space.
@@
From properties (35a-c), one can show that
@@
[…]
From properties (35b,c), it follows that  [^4] Property (35b) can be re-written as $(\mathcal{A}^{\mathcal{D}})^{2}\mathcal{A}=\mathcal{A}^{\mathcal{D}}$ by the commutative property (35c). Thus $\mathcal{A}x=0\implies\mathcal{A}^{\mathcal{D}}x=0$. An identical argument holds for right-multiplication.
@@ V.2 Operator decomposition
Lastly, since a protocol $V(t)$ can be completely characterized up to an (irrelevant) additive constant by the evolution of the equilibrium distribution $\pi(t)$, we use the operator decomposition to derive the relation (App. E)
@@ V.3 Connection to integral relaxation time
As noted in [11], the friction tensor $\tilde{\zeta}$ (for any set of control parameters) may be written as
@@ V.4 Spectral decomposition
The zeroth conjugate relaxation mode $\tilde{\bm{\psi}}_{0}$ is the equilibrium average $-\left\langle\bm{f}\right\rangle_{\pi}$ of the conjugate-force vector, and for $\alpha\geq 1$, $\tilde{\bm{\psi}}_{\alpha}$ captures the fixed-parameter relaxation of $-\left\langle\delta\bm{f}(t)\right\rangle$ corresponding to system relaxation along mode $\alpha$. The linear-response excess power in terms of a limited set of control parameters is then
@@
Equation (57) is particularly powerful for continuous systems for which the spectrum of $\mathcal{L}$ is analytically known. While the full-control spectral decomposition (54) is a manifestly infinite sum for continuous state spaces, the partial-control spectral decomposition (57) may be a finite sum if the conjugate forces $f_{i}(x)$ are orthogonal to all but a finite set of modes $\psi_{\alpha}(x)$; Section (VI.2) demonstrates a concrete example of this.
VI Illustrative examples
Figure 4: Model systems studied in Sec. VI: (a) two-state system with transition rates $w_{01}$ and $w_{10}$ (left), one instantiation of which is a quantum dot interacting with a metallic reservoir [5] (right); (b) overdamped diffusion $X_{t}$ of a particle in a harmonic trap $V(x,t)$ parameterized by tunable stiffness $\kappa(t)$ and center $x_{0}(t)$; (c) four-spin system with single spin-flip dynamics, giving rise to a state graph that is isomorphic to a hypercube (left). Control-parameter sets are (possibly constrained) collections of $k$-spin couplings ($k=1,2,3,4$) drawn from 15 parameters of the cluster expansion (72), e.g., the 5-parameter set $(h,J_{\text{NN}},J_{\text{NNN}},K,L)$ where $h=h_{0}=h_{1}=h_{2}=h_{3}$, $J_{\text{NN}}=J_{01}=J_{02}=J_{13}=J_{23}$, $J_{\text{NNN}}=J_{03}=J_{12}$, and $K=K_{012}=K_{013}=K_{023}=K_{123}$ (right).
VI.1 Two-state system
We begin with a very simple model system: two states with arbitrary transition rates $w_{10}$ and $w_{01}$ (Fig. 4a, left). From Eq. (36), the Drazin inverse of the rate matrix is
$\mathbb{W}^{\mathcal{D}}=\frac{1}{(w_{01}+w_{10})^{2}}\begin{bmatrix}-w_{10}&w_{01}\\ w_{10}&-w_{01}\end{bmatrix}\ .$ (60)
@@ VI.1 Two-state system
$\zeta=\beta\,\pi_{0}(t)\,\pi_{1}(t)\,\tau_{\text{rel}}(t)\begin{bmatrix}1&-1\\ -1&1\end{bmatrix}\ ,$ (61)
@@
A physical instantiation of a two-state system is a single-level quantum dot interacting with a metallic reservoir (Fig. 4a, right). In the wide-band approximation (i.e., the metallic reservoir’s density of states is assumed constant [50]), the energy $\varepsilon$ of the quantum dot and the chemical potential $\mu$ of the reservoir set the respective tunneling rates $w_{10}=2R_{0}[1+\text{e}^{\beta(\mu-\varepsilon)}]^{-1}$ into the reservoir and $w_{01}=2R_{0}[1+\text{e}^{\beta(\varepsilon-\mu)}]^{-1}$ out of the reservoir, where $R_{0}$ is the tunneling rate when $\mu=\varepsilon$ [5]. Then the friction tensor (for one control parameter, the friction coefficient) is
@@
VI.2 Harmonic trap
Consider the overdamped diffusion $X_{t}$ of a particle in a one-dimensional harmonic trap $V(x,t)=\frac{1}{2}\kappa(t)(x-x_{0}(t))^{2}$ with stiffness $\kappa(t)$ and trap center $x_{0}(t)$ (Fig. 4b). The Fokker-Planck operator governing the evolution of the particle’s position distribution $p(x,t)$ is
@@ VI.2 Harmonic trap
$\tilde{\bm{\psi}}_{2}$ $=\frac{1}{\sqrt{2}\kappa\beta}\begin{pmatrix}0\\ 1\end{pmatrix}\ ,$ (70b)
@@
$=\begin{pmatrix}\gamma&0\\ 0&\frac{\gamma}{4\beta\kappa^{3}}\end{pmatrix}\ ,$ (71b)
@@
VI.3 Classical spin systems
VI.3.1 Complete control set for $N$ spins
@@
where $\left\{X^{k}\right\}$ is the set of all $n$-spin interactions with $n=1,\dots,N$, and $\mathcal{C}(k)$ is the set of spin indices corresponding to the coupling $X^{k}$. See, e.g., (VI.3.2) for four spins. This energy function is often called the cluster expansion [54], and it arises in the study of spin glasses and machine learning as the generalized Ising model [55], generalized lattice-gas model [56], or generalized restricted Boltzmann machine [57].
@@
Thus, there is an invertible transformation between the full set of spin moments $-\left\langle\bm{f}\right\rangle$ and the probability distribution $\bm{p}$. We may identify $\mathbb{J}^{\mathsf{R}}=2^{-N}\mathbb{J}^{\mathsf{T}}$ from the above expression or derive it directly (App. F). This invertibility means that the dynamics of the full set of central moments $\left\langle\delta f_{k}\right\rangle$ is fully specified by their current values:
@@
VI.3.2 Four-spin model
@@
Figure 5: (a) Linear-response (LR) excess work for LR optimal protocols for the four-spin system computed for 200 different control sets with between $1$ and $15$ control parameters. The protocols featured in Fig. 6 and Fig. 7 are labeled. (b) 2D histogram showing the number of independently varying parameters against the number of available control parameters (e.g., 18 of the calculated 5-parameter optimal protocols had only 4 independently varying parameters).
@@
From these 15 parameters, the imposition of constraints on their time evolution may be used to build partial control sets. Figure 4c shows one such partial control set, with a bulk field $h=h_{0}=h_{1}=h_{2}=h_{3}$, two-spin coupling blocks $J_{\text{NN}}=J_{01}=J_{02}=J_{13}=J_{23}$ (for nearest neighbors) and $J_{\text{NNN}}=J_{03}=J_{12}$ (for next-nearest neighbors), bulk three-spin coupling $K=K_{012}=K_{013}=K_{023}=K_{123}$, and four-spin coupling $L$.
We set our boundary conditions in terms of these 5 bulk couplings to $h(0)=-1,J_{\text{NN}}(0)=1$ and $h(\tau)=+1,J_{\text{NN}}(\tau)=1$, with all other couplings set to zero at $t=0$ and $t=\tau$. Though (VI.3.2) provides no notion of “geometry” for the spin system, under these boundary conditions it is constructive to imagine a ferromagnetic Ising model on a $2\times 2$ grid as in Fig. 4c.
Figure 6: (a) Linear-response (LR) optimal control protocols for a classical four-spin system in scaled time $s=t/\tau$, for three different control-parameter sets. (b) Spectral contributions to the LR excess power (54). The dissipation along mode $\alpha$ is the product of the relaxation time $\tau_{\alpha}(s)$ and the squared inner product $\left(\mathrm{d}\beta V/\mathrm{d}s,\psi_{\alpha}\right)^{2}$ of the control-parameter velocity with the instantaneous mode. Modes are ordered such that $\tau_{1}(0)\geq\tau_{2}(0)\geq\cdots$. In the middle panel, the squared inner products of the degenerate modes $7$ and $8$ with the velocity are summed since their respective contributions to the dissipation cannot meaningfully be distinguished. Dotted black lines: modes that are everywhere orthogonal (or nearly so) to the velocity $\mathrm{d}\beta V/\mathrm{d}s$ and which thus do not meaningfully contribute to dissipation.
@@
The results are summarized in Fig. 5. We observe that very few protocols need to use all available degrees of freedom to minimize the linear-response excess work. For instance, the optimized 15-parameter full-control protocol is (to numerical precision) identical to the 5-parameter protocol (Fig. 5a blue diamond, and Fig. 6 right). Evidently (at least within the linear-response regime) it is not possible to reduce dissipation by assuming a finer level of control over the spin system. In all, only 42 distinct protocols were identified among the 200 computed. Figure 5b shows the number of independently varying parameters compared to the number of available parameters; at most 7 parameters were utilized.
Figure 6 shows LR-optimized protocols for 3 different sets of control parameters: the one-parameter set $(h,\neg J,\neg K,\neg L)$, the three-parameter set $(h_{=},\neg J,K,\neg L)$ and the five-parameter set $(h,(J_{\text{NN}},J_{\text{NNN}}),K,L)$. Here the four elements of the sets refer to constraints on control of the $1$-, $2$-, $3$-, and $4$-spin parameters, respectively (see Table 1 of App. G). For instance, $h_{=}$ represents two wide magnetic fields, one interacting with the upper row of spins ($\sigma_{0},\sigma_{1}$) and the other with the bottom row ($\sigma_{2},\sigma_{3}$), realized by constraints $h_{0}=h_{1}$ and $h_{2}=h_{3}$. The symbol $\neg$ indicates no control of the given type of coupling, so $\neg J$ indicates that $J_{\text{NN}}=1$ and $J_{\text{NNN}}=0$ are fixed.
@@
The other two protocols (Fig. 6a, middle and right) are geodesics on three- and five-dimensional thermodynamic manifolds. Both protocols change the magnetic fields nonmonotonically, a feature previously observed in [34]. How this (perhaps surprising) behavior reduces the total excess work of the protocols is revealed by examining the separate spectral contributions to the LR excess power (55).
@@
For the 3-parameter protocol, not only is $\tau_{1}$ not significantly reduced, it is roughly doubled near the ends of the protocol. In breaking the symmetry of the four spins, the driving is dominantly directed along the relatively fast-relaxing second mode $\bm{\psi}_{2}$, which can be interpreted as relaxation between the ferromagnetic states and the state $\bm{\sigma}=(+1,+1,-1,-1)$. We also observe that (somewhat counterintuitively) most of the $\sim$15of dissipation attributable to the first mode occurs near $s=1/2$ (when the spectral gap is relatively small), while most of the $\sim$80of dissipation attributable to the second mode occurs when the spectral gap is largest. While $\tau_{1}$ is often regarded as the characteristic dynamical times for a system with a large spectral gap [62, 63], our results show that in driven systems faster modes can become salient, and may even dominate the dissipation – especially for efficient protocols which may drive orthogonal to slow modes to save energy.
Lastly, Fig. 7 shows the linear-response minimum-work protocol for the two-parameter control set $(h,\neg J,\neg K,L)$. The control strategy is similar to the 5-parameter MWP: in the relevant region of the $(h,L)$-manifold, the projection $\tilde{\bm{\psi}}_{1}$ of the first mode (black arrows) is nearly parallel to the $h$-direction. So, given that one must drive along the first mode in order to flip the spins, reduced dissipation is achieved by driving the system nearly perpendicularly to this mode to reach a region of control-parameter space where $\tau_{1}$ is smaller.
@@
VII Summary and Outlook
We have presented a new foundation for the friction-tensor formalism for conservative control of stochastic systems. Thermodynamic manifolds $\tilde{\mathcal{M}}$ for partial control sets $\bm{u}$ are derived as submanifolds of a global thermodynamic manifold $\mathcal{M}$ whose metric structure is provided by the full-control friction tensor. We derived two decompositions of the full-control friction tensor which facilitate both interpretation and calculation of arbitrary friction tensors and minimum-work protocols.

### v3

@@ V.1 The Drazin inverse
Intuitively, the need for an appropriately defined inverse of the generator arises due to the time integral in the definition (45) of the full-control friction tensor. In the same way as the generator $\mathcal{G}_{t}$ plays the role of a time derivative on the space of probability distributions, its Drazin inverse $\mathcal{G}^{\mathcal{D}}_{t}$ plays the role of a time integral (over $0,\infty)$) on the same space.
@@
From properties (36a-c), one can show that
@@
so that $\mathbb{W}^{\mathcal{D}}$ is a proper inverse of $\mathbb{W}$ when the domain is restricted to vectors in the hyperplane $\left\{\bm{v}\ :\ \bm{1}^{T}\bm{v}=0\right\}$ – including deviations $\delta\bm{p}$ from equilibrium and derivatives $\mathrm{d}\bm{p}/\mathrm{d}t$ of probability distributions.
From properties (36b,c), it follows that  [^4] Property (36b) can be re-written as $(\mathcal{A}^{\mathcal{D}})^{2}\mathcal{A}=\mathcal{A}^{\mathcal{D}}$ by the commutative property (36c). Thus $\mathcal{A}x=0\implies\mathcal{A}^{\mathcal{D}}x=0$. An identical argument holds for right-multiplication.
@@ V.2 Operator decomposition
Finally, since a protocol $V(t)$ can be completely characterized up to an (irrelevant) additive constant by the evolution of the equilibrium distribution $\pi(t)$, we use the operator decomposition to derive the relation (App. E)
@@ V.3 Connection to integral relaxation time
As noted in Ref. [15], the friction tensor $\tilde{\zeta}$ (for any set of control parameters) may be written as
@@ V.4 Spectral decomposition
The zeroth conjugate relaxation mode $\tilde{\bm{\psi}}_{0}$ is the equilibrium average $\left\langle\bm{f}\right\rangle_{\pi}$ of the conjugate-force vector, and for $\alpha\geq 1$, $\tilde{\bm{\psi}}_{\alpha}$ captures the fixed-parameter relaxation of $\left\langle\delta\bm{f}(t)\right\rangle$ corresponding to system relaxation along mode $\alpha$. The linear-response excess power in terms of a limited set of control parameters is then
@@
Equation (58) is particularly powerful for continuous systems for which the spectrum of $\mathcal{L}$ is analytically known. While the full-control spectral decomposition (55) is a manifestly infinite sum for continuous state spaces, the partial-control spectral decomposition (58) may be a finite sum if the conjugate forces $f_{i}(x)$ are orthogonal to all but a finite set of modes $\psi_{\alpha}(x)$; Section (VII.2) demonstrates a concrete example of this.
VI Theoretical synopsis
We have, in the preceding sections, presented a rich geometric structure of dissipation under slow driving. Here we briefly review the essential features of this structure before applying it in three illustrative examples.
To each dynamical system evolving reversibly and ergodically under the dynamics (4), one may define a global thermodynamic manifold whose intrinsic geometry captures the relaxation dynamics, as made evident by the decomposition of the metric tensor in terms of the spectrum of the infinitesimal generator (55). This geometry also captures dissipation for slow driving: the squared length of the tangent vector approximates the excess power dissipation for a particular speed and direction of driving (26). This generalizes the friction tensor defined in [15], and subsumes parameter-dependent friction tensors as inherited metrics (30) on submanifolds. The operator decomposition (43) of the full-control metric identifies an object that has implicitly arisen in different forms in previous works [46, 45, 27], unifying its discrete- and continuous-state presentations.
We next demonstrate in examples the computational advantages conferred by operator decomposition and the submanifold structure, provide a concrete example of a Riemannian full-control manifold, and explore the analytical tools offered by the full- and partial-control spectral decompositions.
VII Illustrative examples
Figure 4: Model systems studied in Sec. VII: (a) two-state system with transition rates $w_{01}$ and $w_{10}$ (left), one instantiation of which is a quantum dot interacting with a metallic reservoir [8] (right); (b) overdamped diffusion $X_{t}$ of a particle in a harmonic trap $V(x,t)$ parameterized by tunable stiffness $\kappa(t)$ and center $x_{0}(t)$; (c) four-spin system with single spin-flip dynamics, giving rise to a state graph that is isomorphic to a hypercube (left). Control-parameter sets are (possibly constrained) collections of $k$-spin couplings ($k=1,2,3,4$) drawn from 15 parameters of the cluster expansion (73), e.g., the 5-parameter set $(h,J_{\text{NN}},J_{\text{NNN}},K,L)$ where $h=h_{0}=h_{1}=h_{2}=h_{3}$, $J_{\text{NN}}=J_{01}=J_{02}=J_{13}=J_{23}$, $J_{\text{NNN}}=J_{03}=J_{12}$, and $K=K_{012}=K_{013}=K_{023}=K_{123}$ (right).
VII.1 Two-state system
We begin with a very simple model system: two states with arbitrary transition rates $w_{10}$ and $w_{01}$ [Fig. 4(a), left]. From Eq. (37), the Drazin inverse of the rate matrix is
$\mathbb{W}^{\mathcal{D}}=\frac{1}{(w_{01}+w_{10})^{2}}\begin{bmatrix}-w_{10}&w_{01}\\ w_{10}&-w_{01}\end{bmatrix}\$. (61)
@@ VII.1 Two-state system
$\zeta=\beta\,\pi_{0}(t)\,\pi_{1}(t)\,\tau_{\text{rel}}(t)\begin{bmatrix}1&-1\\ -1&1\end{bmatrix}\$, (62)
@@
A physical instantiation of a two-state system is a single-level quantum dot interacting with a metallic reservoir [Fig. 4(a), right]. In the wide-band approximation (i.e., the metallic reservoir’s density of states is assumed constant [55]), the energy $\varepsilon$ of the quantum dot and the chemical potential $\mu$ of the reservoir set the respective tunneling rates $w_{10}=2R_{0}[1+\text{e}^{\beta(\mu-\varepsilon)}]^{-1}$ into the reservoir and $w_{01}=2R_{0}[1+\text{e}^{\beta(\varepsilon-\mu)}]^{-1}$ out of the reservoir, where $R_{0}$ is the tunneling rate when $\mu=\varepsilon$ [8]. Then the friction tensor (for one control parameter, the friction coefficient) is
@@
VII.2 Harmonic trap
Consider the overdamped diffusion $X_{t}$ of a particle in a one-dimensional harmonic trap $V(x,t)=\frac{1}{2}\kappa(t)(x-x_{0}(t))^{2}$ with stiffness $\kappa(t)$ and trap center $x_{0}(t)$ [Fig. 4(b)]. The Fokker-Planck operator governing the evolution of the particle’s position distribution $p(x,t)$ is
@@ VII.2 Harmonic trap
$\tilde{\bm{\psi}}_{2}$ $=\frac{1}{\sqrt{2}\kappa\beta}\begin{pmatrix}0\\ 1\end{pmatrix}\$, (71b)
@@
$=\begin{pmatrix}\gamma&0\\ 0&\frac{\gamma}{4\beta\kappa^{3}}\end{pmatrix}\$, (72b)
@@
VII.3 Classical spin systems
VII.3.1 Complete control set for $N$ spins
@@
where $\left\{X^{k}\right\}$ is the set of all $n$-spin interactions with $n=1,\dots,N$, and $\mathcal{C}(k)$ is the set of spin indices corresponding to the coupling $X^{k}$. See, e.g., (VII.3.2) for four spins. This energy function is often called the cluster expansion [58], and it arises in the study of spin glasses and machine learning as the generalized Ising model [59], generalized lattice-gas model [60], or generalized restricted Boltzmann machine [61].
@@
Thus, there is an invertible transformation between the full set of spin moments $\left\langle\bm{f}\right\rangle$ and the probability distribution $\bm{p}$. We may identify $\mathbb{J}^{\mathsf{R}}=2^{-N}\mathbb{J}^{\mathsf{T}}$ from the above expression or derive it directly (App. F). This invertibility means that the dynamics of the full set of central moments $\left\langle\delta f_{k}\right\rangle$ is fully specified by their current values:
@@
VII.3.2 Four-spin model
@@
Figure 5: (a) LR excess work for LR optimal protocols for the four-spin system computed for 200 different control sets with between $1$ and 1515 control parameters. The protocols featured in Fig. 6 and Fig. 7 are labeled. (b) Two-dimensional histogram showing the number of independently varying parameters against the number of available control parameters (e.g., 18 of the calculated 5-parameter optimal protocols had only 4 independently varying parameters).
@@
From these 15 parameters, the imposition of constraints on their time evolution may be used to build partial control sets. Figure 4(c) shows one such partial control set, with a bulk field $h=h_{0}=h_{1}=h_{2}=h_{3}$, two-spin coupling blocks $J_{\text{NN}}=J_{01}=J_{02}=J_{13}=J_{23}$ (for nearest neighbors) and $J_{\text{NNN}}=J_{03}=J_{12}$ (for next-nearest neighbors), bulk three-spin coupling $K=K_{012}=K_{013}=K_{023}=K_{123}$, and four-spin coupling $L$.
We set our boundary conditions in terms of these 5 bulk couplings to $h(0)=-1,J_{\text{NN}}(0)=1$ and $h(\tau)=+1,J_{\text{NN}}(\tau)=1$, with all other couplings set to zero at $t=0$ and $t=\tau$. Though (VII.3.2) provides no notion of “geometry” for the spin system, under these boundary conditions it is constructive to imagine a ferromagnetic Ising model on a $2\times 2$ grid as in Fig. 4(c).
Figure 6: (a) LR optimal control protocols for a classical four-spin system in scaled time $s=t/\tau$, for three different control-parameter sets. (b) Spectral contributions to the LR excess power (55). The dissipation along mode $\alpha$ is the product of the relaxation time $\tau_{\alpha}(s)$ and the squared inner product $\left(\mathrm{d}\beta V/\mathrm{d}s,\psi_{\alpha}\right)^{2}$ of the control-parameter velocity with the instantaneous mode. Modes are ordered such that $\tau_{1}(0)\geq\tau_{2}(0)\geq\cdots$. In the middle panel, the squared inner products of the degenerate modes 77 and 88 with the velocity are summed since their respective contributions to the dissipation cannot meaningfully be distinguished. Dotted black lines: modes that are everywhere orthogonal (or nearly so) to the velocity $\mathrm{d}\beta V/\mathrm{d}s$ and which thus do not meaningfully contribute to dissipation.
@@
The results are summarized in Fig. 5. We observe that very few protocols need to use all available degrees of freedom to minimize the linear-response excess work. For instance, the optimized 15-parameter full-control protocol is (to numerical precision) identical to the 5-parameter protocol [Fig. 5(a) blue diamond, and Fig. 6 right]. Evidently (at least within the linear-response regime) it is not possible to reduce dissipation by assuming a finer level of control over the spin system. In all, only 42 distinct protocols were identified among the 200 computed. Figure 5b shows the number of independently varying parameters compared to the number of available parameters; at most 7 parameters were utilized.
Figure 6 shows LR-optimized protocols for 3 different sets of control parameters: the one-parameter set $(h,\neg J,\neg K,\neg L)$, the three-parameter set $(h_{=},\neg J,K,\neg L)$ and the five-parameter set $(h,(J_{\text{NN}},J_{\text{NNN}}),K,L)$. Here the four elements of the sets refer to constraints on control of the $1$-, 22-, 33-, and $4$-spin parameters, respectively (see Table 1 of App. G). For instance, $h_{=}$ represents two wide magnetic fields, one interacting with the upper row of spins ($\sigma_{0},\sigma_{1}$) and the other with the bottom row ($\sigma_{2},\sigma_{3}$), realized by constraints $h_{0}=h_{1}$ and $h_{2}=h_{3}$. The symbol $\neg$ indicates no control of the given type of coupling, so $\neg J$ indicates that $J_{\text{NN}}=1$ and $J_{\text{NNN}}=0$ are fixed.
@@
The other two protocols [Fig. 6(a), middle and right] are geodesics on three- and five-dimensional thermodynamic manifolds. Both protocols change the magnetic fields nonmonotonically, a feature previously observed in Ref. [39]. How this (perhaps surprising) behavior reduces the total excess work of the protocols is revealed by examining the separate spectral contributions to the LR excess power (56).
@@
For the 3-parameter protocol, not only is $\tau_{1}$ not significantly reduced, it is roughly doubled near the ends of the protocol. In breaking the symmetry of the four spins, the driving is dominantly directed along the relatively fast-relaxing second mode $\bm{\psi}_{2}$, which can be interpreted as relaxation between the ferromagnetic states and the state $\bm{\sigma}=(+1,+1,-1,-1)$. We also observe that (somewhat counterintuitively) most of the $\sim$15% of dissipation attributable to the first mode occurs near $s=1/2$ (when the spectral gap is relatively small), while most of the $\sim$80% of dissipation attributable to the second mode occurs when the spectral gap is largest. While $\tau_{1}$ is often regarded as the characteristic dynamical times for a system with a large spectral gap [66, 67], our results show that in driven systems faster modes can become salient, and may even dominate the dissipation – especially for efficient protocols which may drive orthogonal to slow modes to save energy.
Finally, Fig. 7 shows the linear-response minimum-work protocol for the two-parameter control set $(h,\neg J,\neg K,L)$. The control strategy is similar to the 5-parameter MWP: in the relevant region of the $(h,L)$-manifold, the projection $\tilde{\bm{\psi}}_{1}$ of the first mode (black arrows) is nearly parallel to the $h$-direction. So, given that one must drive along the first mode in order to flip the spins, reduced dissipation is achieved by driving the system nearly perpendicularly to this mode to reach a region of control-parameter space where $\tau_{1}$ is smaller.
@@
VIII Summary and Outlook
We have presented a foundation for the friction-tensor formalism for conservative control of stochastic systems. Thermodynamic manifolds $\tilde{\mathcal{M}}$ for partial control sets $\bm{u}$ are derived as submanifolds of a global thermodynamic manifold $\mathcal{M}$ whose metric structure is provided by the full-control friction tensor. We derived two decompositions of the full-control friction tensor which facilitate both interpretation and calculation of arbitrary friction tensors and minimum-work protocols.